# Improving Retrieval-Augmented Generation in Medicine with Iterative Follow-up Questions

### v1

@@ Abstract
The emergent abilities of large language models (LLMs) have demonstrated great potential in solving medical questions. They can possess considerable medical knowledge, but may still hallucinate and are inflexible in the knowledge updates. While Retrieval-Augmented Generation (RAG) has been proposed to enhance the medical question-answering capabilities of LLMs with external knowledge bases, it may still fail in complex cases where multiple rounds of information-seeking are required. To address such an issue, we propose iterative RAG for medicine (i-MedRAG), where LLMs can iteratively ask follow-up queries based on previous information-seeking attempts. In each iteration of i-MedRAG, the follow-up queries will be answered by a vanilla RAG system and they will be further used to guide the query generation in the next iteration. Our experiments show the improved performance of various LLMs brought by i-MedRAG compared with vanilla RAG on complex questions from clinical vignettes in the United States Medical Licensing Examination (USMLE), as well as various knowledge tests in the Massive Multitask Language Understanding (MMLU) dataset. Notably, our zero-shot i-MedRAG outperforms all existing prompt engineering and fine-tuning methods on GPT-3.5, achieving an accuracy of 69.68% on the MedQA dataset. In addition, we characterize the scaling properties of i-MedRAG with different iterations of follow-up queries and different numbers of queries per iteration. Our case studies show that i-MedRAG can flexibly ask follow-up queries to form reasoning chains, providing an in-depth analysis of medical questions. To the best of our knowledge, this is the first-of-its-kind study on incorporating follow-up queries into medical RAG.
@@ 1 Introduction
While there have been several medical applications of RAG, such as Almanac [21], Clinfo.ai [22], and MedRAG [23], their RAG component is mainly beneficial to questions that have direct answers in a single document, such as those in the PubMedQA [24] and BioASQ [25] datasets. However, only marginal improvements are seen with RAG for questions that require multiple rounds of clinical reasoning like MedQA [26], a dataset curated from medical license examinations. For example, to recommend a treatment for a patient with certain symptoms, a system needs to first infer the potential diagnosis from the symptoms and then find a suitable treatment for the diagnosis. Nevertheless, only one round of retrieval is conducted in the vanilla RAG architecture, prohibiting multiple rounds of information seeking that are required in complex clinical reasoning.
In this work, we propose i-MedRAG, a simple and effective framework for incorporating follow-up queries into RAG. Specifically, we prompt LLMs to iteratively generate follow-up queries to search for additional information from external medical corpora. The queries and the corresponding answers generated with RAG will be used to augment the answer generation of the original question. Empirical results demonstrate the effectiveness of i-MedRAG on both open- and close-source LLMs, which show improved performance on the United States Medical Licensing Examination (USMLE) subset of MedQA and medical questions from the Massive Multitask Language Understanding (MMLU) dataset. Notably, our zero-shot i-MedRAG outperforms all previous prompt engineering and fine-tuning methods, setting a state-of-the-art accuracy for GPT-3.5 on the MedQA dataset (Figure 1). Our further analysis of the number of iterations and number of queries per iteration used in i-MedRAG reflects how its performance scales with different settings. Additionally, we present several case studies of i-MedRAG, showing how it overcomes the limitations in vanilla RAG to find the correct answers.
@@
  • We have conducted comprehensive experiments on medical question answering, and the results demonstrate that i-MedRAG not only outperforms vanilla RAG approaches but also surpasses all other prompt engineering approaches on MedQA with GPT-3.5, setting a new state-of-the-art performance of 69.68%.
@@ 2.1 Retrieval-Augmented Generation for Medicine
Retrieval-augmented generation (RAG) has been widely adopted in medicine. Here, we discuss several representative approaches. Almanac [21] is a system that augments LLMs with curated resources for medical guidelines and treatment recommendations, which shows improvements over the standard LLMs in six manually assessed metrics. Similarly, Low et al. [27] demonstrate the improvements of RAG-based systems for real-world clinical queries with manual evaluation. Clinfo.ai [22] is an open-source web application that answers clinical questions based on retrieved scientific literature from PubMed articles. Xiong et al. [23] conduct a benchmarking study with the MedRAG toolkit, and show the benefits of RAG in several medical multi-choice question answering datasets. There are also various biomedical literature search products [28] that use RAG to summarize the retrieved articles [29], such as OpenEvidence [^1] https://www.openevidence.com/ and ChatRWD [^2] https://www.atroposhealth.com/chatrwd. However, most of the RAG studies in medicine use the vanilla architecture with only one round of retrieval. To the best of our knowledge, our study presents the first approach and evaluations on incorporating follow-up queries in RAG for medicine.
@@ 3 Methods
Figure 2 shows the overview of our i-MedRAG and its comparison to the vanilla Retrieval-Augmented Generation (RAG). Different from RAG, our i-MedRAG modifies its pipeline by replacing the information retrieval step (Figure 2 left) with our proposed iterative question-answering step (Figure 2 middle and right). The settings of RAG are described in Section 3.1 and the pipeline of our new i-MedRAG is discussed in Section 3.2. The details of the iterative question answering are described in Section 3.3.

### v2

@@ Abstract
The emergent abilities of large language models (LLMs) have demonstrated great potential in solving medical questions. They can possess considerable medical knowledge, but may still hallucinate and are inflexible in the knowledge updates. While Retrieval-Augmented Generation (RAG) has been proposed to enhance the medical question-answering capabilities of LLMs with external knowledge bases, it may still fail in complex cases where multiple rounds of information-seeking are required. To address such an issue, we propose iterative RAG for medicine (i-MedRAG), where LLMs can iteratively ask follow-up queries based on previous information-seeking attempts. In each iteration of i-MedRAG, the follow-up queries will be answered by a conventional RAG system and they will be further used to guide the query generation in the next iteration. Our experiments show the improved performance of various LLMs brought by i-MedRAG compared with conventional RAG on complex questions from clinical vignettes in the United States Medical Licensing Examination (USMLE), as well as various knowledge tests in the Massive Multitask Language Understanding (MMLU) dataset. Notably, our zero-shot i-MedRAG outperforms all existing prompt engineering and fine-tuning methods on GPT-3.5, achieving an accuracy of 69.68% on the MedQA dataset. In addition, we characterize the scaling properties of i-MedRAG with different iterations of follow-up queries and different numbers of queries per iteration. Our case studies show that i-MedRAG can flexibly ask follow-up queries to form reasoning chains, providing an in-depth analysis of medical questions. To the best of our knowledge, this is the first-of-its-kind study on incorporating follow-up queries into medical RAG.
@@ 1 Introduction
While there have been several medical applications of RAG, such as Almanac [21], Clinfo.ai [22], and MedRAG [23], their RAG component is mainly beneficial to questions that have direct answers in a single document, such as those in the PubMedQA [24] and BioASQ [25] datasets. However, only marginal improvements are seen with RAG for questions that require multiple rounds of clinical reasoning like MedQA [26], a dataset curated from medical license examinations. For example, to recommend a treatment for a patient with certain symptoms, a system needs to first infer the potential diagnosis from the symptoms and then find a suitable treatment for the diagnosis. Nevertheless, only one round of retrieval is conducted in the conventional RAG architecture, prohibiting multiple rounds of information seeking that are required in complex clinical reasoning.
In this work, we propose i-MedRAG, a simple and effective framework for incorporating follow-up queries into RAG. Specifically, we prompt LLMs to iteratively generate follow-up queries to search for additional information from external medical corpora. The queries and the corresponding answers generated with RAG will be used to augment the answer generation of the original question. Empirical results demonstrate the effectiveness of i-MedRAG on both open- and close-source LLMs, which show improved performance on the United States Medical Licensing Examination (USMLE) subset of MedQA and medical questions from the Massive Multitask Language Understanding (MMLU) dataset. Notably, our zero-shot i-MedRAG outperforms all previous prompt engineering and fine-tuning methods, setting a state-of-the-art accuracy for GPT-3.5 on the MedQA dataset (Figure 1). Our further analysis of the number of iterations and number of queries per iteration used in i-MedRAG reflects how its performance scales with different settings. Additionally, we present several case studies of i-MedRAG, showing how it overcomes the limitations in conventional RAG to find the correct answers.
@@
  • We have conducted comprehensive experiments on medical question answering, and the results demonstrate that i-MedRAG not only outperforms conventional RAG approaches but also surpasses all other prompt engineering approaches on MedQA with GPT-3.5, setting a new state-of-the-art performance of 69.68%.
@@ 2.1 Retrieval-Augmented Generation for Medicine
Retrieval-augmented generation (RAG) has been widely adopted in medicine. Here, we discuss several representative approaches. Almanac [21] is a system that augments LLMs with curated resources for medical guidelines and treatment recommendations, which shows improvements over the standard LLMs in six manually assessed metrics. Similarly, Low et al. [27] demonstrate the improvements of RAG-based systems for real-world clinical queries with manual evaluation. Clinfo.ai [22] is an open-source web application that answers clinical questions based on retrieved scientific literature from PubMed articles. Xiong et al. [23] conduct a benchmarking study with the MedRAG toolkit, and show the benefits of RAG in several medical multi-choice question answering datasets. There are also various biomedical literature search products [28] that use RAG to summarize the retrieved articles [29], such as OpenEvidence [^1] https://www.openevidence.com/ and ChatRWD [^2] https://www.atroposhealth.com/chatrwd. However, most of the RAG studies in medicine use the conventional architecture with only one round of retrieval. To the best of our knowledge, our study presents the first approach and evaluations on incorporating follow-up queries in RAG for medicine.
@@ 3 Methods
Figure 2 shows the overview of our i-MedRAG and its comparison to the conventional Retrieval-Augmented Generation (RAG). Different from RAG, our i-MedRAG modifies its pipeline by replacing the information retrieval step (Figure 2 left) with our proposed iterative question-answering step (Figure 2 middle and right). The settings of RAG are described in Section 3.1 and the pipeline of our new i-MedRAG is discussed in Section 3.2. The details of the iterative question answering are described in Section 3.3.